# Fully Parallelized BP Decoding for Quantum LDPC Codes Can Outperform BP-OSD

### v2

@@ 3.2. Oscillation
As suggested in previous works (Raveendran and Vasić, 2021; Chytas et al., 2024), trapping sets and code degeneracy often result in ambiguous BP decoding, leading to convergence failures. A common symptom of such a failure is bit-level oscillation, where certain output bits repeatedly flip between 0 and 1 across iterations. To better understand the relationship between oscillating bits and decoding errors, we analyze the dynamics of bit oscillations during the BP process. During the decoding process, we track bit-level oscillations by comparing the output of each iteration with that of the previous one and counting how often each bit flips, which is similar to Ref (Chytas et al., 2024). We then identify a set of oscillating bits, denoted by $\Phi$, based on their flip frequency. Specifically, $\Phi$ is defined as the top $|\Phi|$ of the most frequently flipped bits. We denote $\operatorname{supp}(\bm{e})$ as the set of erroneous bits and define the hit precision and recall as
(9) Precision $=\frac{|\operatorname{supp}(\bm{e})\cap\Phi|}{|\Phi|},$
(10) Recall $=\frac{|\operatorname{supp}(\bm{e})\cap\Phi|}{|\operatorname{supp}(\bm{e})|}.$
@@
Fig. 2 shows the precision and recall rate when the min-sum decoder fails to decode a syndrome. We can see that even when the BP decoder fails to fully correct an error, the pattern of bit oscillations essentially reveals a meaningful subset of the actual error locations. In particular, at lower physical error rates, the set of oscillating bits nearly covers the entire true error positions. To confirm, we observe that the hit precision, i.e., the fraction of oscillating bits that are indeed erroneous, is substantially higher than the physical error rate, indicating that this method significantly outperforms random guessing. Therefore, this information can serve as a valuable indicator for identifying unreliable bits during post-processing. As the physical error rate increases, the recall decreases, primarily because the total number of errors grows while the candidate set size remains fixed.
4. Proposed Speculative Decoding Method
@@
To enhance the performance of BP decoding, we adopt a Chase-like post-processing technique (Chase, 1972). This approach generates a set of test vectors by flipping candidate bits and attempts decoding on each of them, thereby increasing the likelihood of successful error correction. In the quantum decoding setting, we only have access to the syndrome rather than the prior probability of each bit being in error. Therefore, unlike the original Chase algorithm, which relies on prior channel information to select candidate bits, our method identifies candidate bits based on BP flipping statistics. As shown in Figure 3, once these candidate bits $\Phi$ are identified, we generate diverse decoding attempts by flipping the input syndrome accordingly across multiple BP instances. This strategy increases the variety in the decoder’s inputs and distinguishes our approach from that in (Koutsioumpas et al., 2025), which does not modify the input syndrome. If the decoder successfully converges on this modified input, we then flip these bits back in the output. This restoration ensures that the final output error matches the original syndrome preserving the validity of the decoding result. Since these candidate bits are likely to correspond to actual error locations, flipping them can effectively reduce the number of errors in the input and equivalently lowers the physical error rate. This reduction not only increases the likelihood of successful decoding but also reduces the number of iterations needed for BP to converge. The detailed pseudo code can be found in Algorithm 1.
@@
Figure 3. A simplified flowchart of the proposed decoder. The full procedure is described in Algorithm 1.
@@
             /* Speculative decoding using test patterns based on $\Phi$/
@@
Function BP_with_oscillation($\bm{s},|\Phi|$):
[…]
@@
[…]
@@
                   return True, $\hat{\bm{e}},\emptyset$
@@
Algorithm 1 BP decoding with Chase-like post-processing
@@
Note that this concept is also widely adopted in classical decoding frameworks (Chase, 1972; Halford and Chugg, 2006; Dimnik and Be’ery, 2009; Wang et al., 2023). In such approaches, multiple modified inputs are generated and decoded independently, and the best result is selected according to the maximum likelihood criterion, typically based on minimum weight or log-likelihood score. Since each decoding attempt is independent, they can be executed in parallel, introducing minimal latency overhead. However, in the quantum setting, the structure of qLDPC codes introduces unique advantages. Due to the degeneracy and high distance of qLDPC codes, and the tendency of BP decoders to favor minimum-weight solutions, successful BP convergence rarely results in a logical error. This is partly because the classical codes $\mathcal{C}_{X}$ and $\mathcal{C}_{Z}$, defined by the parity-check matrices $H_{X}$ and $H_{Z}$, have low minimum distances (recall that the minimum distance of $\mathcal{C}_{X}$ is upper bounded by the row weight of $H_{Z}$, and vice versa). Even if BP converges to a non-optimal codeword within $\mathcal{C}_{X}$ or $\mathcal{C}_{Z}$, the resulting error is likely to be close to the minimal solution in terms of Hamming distance.
When considering the full quantum code defined by both $H_{X}$ and $H_{Z}$, the code distance is significantly higher than distances of $\mathcal{C}_{X}$ and $\mathcal{C}_{Z}$. This motivates our use of a speculative decoding strategy. In our approach, we omit the maximum likelihood selection step: Because of code degeneracy, any solution that satisfies the syndrome is likely to belong to the correct coset, particularly in the low-error regime and for high-distance codes. As a result, we simply return the first valid codeword that satisfies the syndrome among the parallel decoding attempts. This strategy reduces latency while preserving decoding performance.
5. Simulation Results
@@
In this section, we evaluate the performance of the proposed decoder and compare it against the BP-OSD decoder. All BP decoders use the min-sum algorithm with an adaptive damping factor $\alpha=1-2^{i}$, where $i$ is the current number of iterations. The OSD method is OSD-CS in (Roffe et al., 2020), and for briefness, we use labels such as “BP1000-OSD10” to denote a decoder using BP with a maximum of 1,000 iterations followed by OSD-CS of order 10. Each data point is obtained by collecting at least 100 logical errors unless specified otherwise. As a result, the statistical uncertainty is sufficiently small, and error bars are omitted for clarity.
5.1. Code Capacity Model
@@
In the code capacity error model, we decode using all test vectors of weight up to 1, which proves sufficient to achieve satisfactory logical error rates. In many of the codes we tested, the BP decoder already performs well under the code-capacity model, leaving limited room for improvement. However, there exist some exceptions. One such case is the $\llbracket 154,6,16\rrbracket$ coprime BB code introduced in (Wang and Mueller, 2024), where the min-sum decoder performs poorly despite the code’s high distance.
@@
Fig. 4 shows the logical error rates of different decoders on the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the code-capacity noise model. For the proposed decoder, the candidate set size is set to $|\Phi|=8$, resulting in a maximum of $50\times(8+1)=450$ BP iterations per decoding attempt. Considering that BP decoders can be run in parallel after the initial run, the latency is equivalent to 100 BP iterations. As shown, our proposed decoder significantly outperforms both the baseline BP and BP-OSD decoders, achieving lower logical error rates with fewer iterations and without requiring costly OSD post-processing. Additionally, both BP and BP-OSD exhibit an error floor at low physical error rates. Upon examining the decoding failures, we find that many are due to low-weight (e.g., weight-3) errors that were caused by trapping sets.
@@
Figure 4. Error rates of the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the code capacity model.
@@
Another example is the $\llbracket 288,12,18\rrbracket$ BB code from (Bravyi et al., 2024). As shown in Fig. 5, the proposed decoder performs similar to the BP-OSD decoder while using fewer than $50\times(20+1)=1050$ iterations per decoding attempt and the latency is still 100 iterations, as we analyzed above.
@@
Figure 5. Error rates of the $\llbracket 288,12,18\rrbracket$ BB code under the code capacity model.
5.2. Circuit-Level Noise Model
@@
For the circuit-level noise model, errors are injected with uniform probability for gates and measurements. We use Stim (Gidney, 2021) to generate the detector error model and the corresponding parity-check matrix, where each row represents a detector event and each column corresponds to an error mechanism. Following the convention in prior literature, we perform $d$ rounds of syndrome extraction and define the logical error rate per round as
(11) LER Per $\text{LER Per Round}=1-(1-\text{LER})^{\frac{1}{d}},$
@@
In this model, the number of error mechanisms is typically much larger than the number of qubits, resulting in very large parity-check matrices. Consequently, flipping a single bit as in the code capacity model is often insufficient for the BP decoder to converge. On the other hand, exhaustively decoding all test vectors with weight up to a threshold is computationally expensive. To address this, we adopt a sampling-based approach. Given a maximum test vector weight $w_{max}$, we randomly sample $n_{s}$ test vectors for each weight in $\{1,\dots,w_{max}\}$, resulting in a total of $n_{s}\times w_{max}$ test vectors per failed BP decoding attempt.
@@
Figure 6 shows the logical error rates of different decoders on the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model. In this setting, the number of error mechanisms is approximately 8,000, which is significantly larger than the code length used in the code-capacity model. Therefore, we expand the candidate set size in the proposed decoder. Specifically, the proposed decoder uses up to 3,100 BP iterations (based on $w_{max}=6,n_{s}=5$) and achieves logical error rates that are slightly higher but still comparable to that of the BP-OSD decoder, which uses a maximum of 1,000 iterations and OSD order 10.
@@
Figure 6. Error rates of the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model.
@@
Figure 7. Error rates of the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the circuit-level noise model.
@@
Figure 7 shows the logical error rates of different decoders on the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model. The proposed decoder uses up to 6,000 iterations (based on $w_{max}=6,n_{s}=10$. However, since all test vectors are decoded in parallel, the effective decoding latency corresponds to only 200 iterations. The proposed decoder achieves logical error rates that are slightly higher but still comparable to that of the BP-OSD decoder at lower physical error rate, which uses a maximum of 1,000 iterations and OSD order 10. In the higher physical error rate regime, the proposed decoder exhibits logical error rates that are higher than those of BP-OSD but still consistently lower than baseline BP decoding.
@@
Figure 8. Error rates of the $\llbracket 126,12,10\rrbracket$ coprime-BB code under the circuit-level noise model.
@@
Figure 8 shows the logical error rates of different decoders on the $\llbracket 126,12,10\rrbracket$ coprime-BB code under the circuit-level noise model. The proposed decoder uses up to approximately 3,000 BP iterations to achieve a logical error rate comparable to that of the BP1000-OSD10 decoder. By increasing both $n_{s}$ and $w_{\max}$, we are able to further reduce the logical error rate to slightly below that of the BP-OSD decoder. However, this improvement comes at the cost of increased complexity, requiring up to 10,000 BP iterations.
5.3. Complexity and Parameter Selection

### v3

@@ III-B Oscillation
As suggested in previous works [20, 5], trapping sets and code degeneracy often result in ambiguous BP decoding, leading to convergence failures. A common symptom of such a failure is bit-level oscillation, where certain output bits repeatedly flip between 0 and 1 across iterations. To better understand the relationship between oscillating bits and decoding errors, we analyze the dynamics of bit oscillations during the BP process. During the decoding process, we track bit-level oscillations by comparing the output of each iteration with that of the previous one and counting how often each bit flips, which is similar to [5]. We then identify a set of oscillating bits, denoted by $\Phi$, based on their flip frequency. Specifically, $\Phi$ is defined as the top $|\Phi|$ of the most frequently flipped bits. We denote $\operatorname{supp}(\bm{e})$ as the set of erroneous bits and define the hit precision and recall as
Precision $=\frac{|\operatorname{supp}(\bm{e})\cap\Phi|}{|\Phi|},$ (9)
Recall $=\frac{|\operatorname{supp}(\bm{e})\cap\Phi|}{|\operatorname{supp}(\bm{e})|}.$ (10)
@@
Fig. 3 shows the precision and recall rate when the min-sum decoder fails to decode a syndrome. We can see that even when the BP decoder fails to fully correct an error, the pattern of bit oscillations essentially reveals a meaningful subset of the actual error locations. In particular, at lower physical error rates, the set of oscillating bits nearly covers the entire true error positions. To confirm, we observe that the hit precision, i.e., the fraction of oscillating bits that are indeed erroneous, is substantially higher than the physical error rate. This suggests that bits in the oscillation set $\Phi$ are much more likely to be true errors than random guesses, making them valuable targets for post-processing. As the physical error rate increases, the recall decreases, primarily because the total number of errors grows while the candidate set size remains fixed.
Figure 3: Precision and recall probabilities of candidate bit selection on the $\llbracket 144,12,12\rrbracket$ code. We evaluate the correlation between candidate bits and actual error locations by identifying the top 50 most frequently flipped bits among approximately 8,000 error mechanisms. The decoder is run with a maximum of 50 iterations, and statistics are collected over 1,000 decoding failures. This analysis reveals how well bit-level oscillation can serve as a heuristic for error localization.
IV A New Speculative Decoding Method
@@
To enhance the performance of BP decoding, we adopt a Chase-like post-processing technique [3] that we call BP-SF (Syndrome Flip). This approach generates a set of trial vectors by flipping candidate bits and attempts decoding on each of them, thereby increasing the likelihood of successful error correction. In the quantum decoding setting, we only have access to the syndrome and prior information of error mechanisms, rather than the prior information of each bit based on signal strength in classical soft decoding. Therefore, unlike the original Chase algorithm, which relies on prior channel information to select candidate bits, our BP-SF method identifies candidate bits based on BP flipping statistics. As shown in Figure 4, once these candidate bits $\Phi$ are identified, we generate diverse decoding attempts by flipping the input syndrome accordingly across multiple BP instances. This strategy increases the variety in the decoder’s inputs and distinguishes our BP-SF approach from that in [15], which modifies the posterior information instead of the syndrome. If the decoder successfully converges on this modified input, we then flip these bits back in the output. This restoration ensures that the final output error matches the original syndrome and preserves the validity of the decoding result. Since these candidate bits are likely to correspond to actual error locations, flipping them can also effectively reduce the number of errors in the input and equivalently lowers the physical error rate. This reduction not only increases the likelihood of successful decoding but also reduces the number of iterations needed for BP to converge. The detailed pseudo code can be found in Algorithm 1.
@@
Figure 4: A simplified flowchart of the proposed decoder. The full procedure is described in Algorithm 1.
@@
    /* Speculative decoding using trial vectors based on $\Phi$ */
@@
0.3em Function BP_with_oscillation($\bm{s},|\Phi|$):
 $\text{flip\_count}\leftarrow\bm{0}$
@@
    $\hat{\bm{e}}\leftarrow\textnormal{{BP\_Update}}()$ /* Standard BP update */
    $\text{flip\_count}\leftarrow\text{flip\_count}+(\hat{\bm{e}}\oplus\hat{\bm{e}}_{\text{prev}})$ /* Track bit oscillations */
@@
       return True, $\hat{\bm{e}}$$\emptyset$
@@
Algorithm 1 BP decoding with our Chase-like post-processing (BP-SF)
@@
We note that this concept is also widely adopted in classical decoding frameworks [3, 12, 9, 32]. In such approaches, multiple modified inputs are generated and decoded independently, and the best result is selected according to the maximum likelihood criterion, typically based on minimum weight or log-likelihood score. Since each decoding attempt is independent, they can be executed in parallel, introducing minimal latency overhead. However, in the quantum setting, the structure of qLDPC codes introduces unique advantages. Due to the degeneracy and high distance of qLDPC codes, and the tendency of BP decoders to favor minimum-weight solutions, successful BP convergence rarely results in a logical error. This is partly because the classical codes $\mathcal{C}_{X}$ and $\mathcal{C}_{Z}$, defined by the parity-check matrices $H_{X}$ and $H_{Z}$, have low minimum distances (recall that the minimum distance of $\mathcal{C}_{X}$ is upper bounded by the row weight of $H_{Z}$, and vice versa). As a result, there are many low-weight codewords in $\mathcal{C}_{X}$ and $\mathcal{C}_{Z}$. Therefore, even if BP converges to a non-optimal codeword within $\mathcal{C}_{X}$ or $\mathcal{C}_{Z}$, the resulting error is likely to differ from the optimal solution by a low-weight codeword. Consequently, the probability that this low-weight codeword forms a logical operator is very low, as the weight of logical operator is much higher (at least $d$). This motivates our use of a speculative decoding strategy. In our BP-SF approach, we omit the maximum likelihood selection step: Because of code degeneracy, any solution that satisfies the syndrome is likely to belong to the correct coset, particularly in the low-error regime and for high-distance codes. As a result, we simply return the first valid codeword that satisfies the syndrome among the parallel decoding attempts. This strategy reduces latency while preserving decoding performance.
V Simulation Results
@@
In this section, we evaluate the performance of our proposed BP-SF decoder and compare it against the BP-OSD decoder. All BP decoders use the min-sum algorithm with an adaptive damping factor $\alpha=1-2^{i}$, where $i$ is the current number of iterations. The OSD method is OSD-CS in [22], and for briefness, we use labels such as “BP1000-OSD10” to denote a decoder using BP with a maximum of 1,000 iterations followed by OSD-CS of order 10. Each data point is obtained by collecting at least 100 logical errors unless specified otherwise. The statistical uncertainty is sufficiently small, and error bars are omitted for clarity.
V-A Code Capacity Model
@@
In the code capacity error model, $X$, $Y$, and $Z$ errors are applied independently for each data qubit with probability of $p/3$. All other operations are perfect, where $p$ is the physical error rate. In this model, we decode all codes using all trial vectors of weight only up to 1, which proves sufficient to achieve satisfactory logical error rates. In many of the codes we tested, the BP decoder already performs well under the code capacity model, leaving limited room for improvement as shown in Appendix -B. However, there exist some exceptions. One such case is the $\llbracket 154,6,16\rrbracket$ coprime BB code introduced in [33], where the min-sum decoder performs poorly despite the code’s high distance.
@@
Fig. 5 shows the logical error rates of different decoders on the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the code capacity noise model. For our BP-SF decoder, the candidate set size is set to $|\Phi|=8$, resulting in a maximum of $50\times(8+1)=450$ BP iterations per decoding attempt. Considering that BP decoders can be run in parallel after the initial run, the latency can be optimized to 100 BP iterations. As shown, our BP-SF decoder significantly outperforms both the baseline BP and BP-OSD decoders, achieving lower logical error rates with fewer iterations and without requiring costly OSD post-processing. Additionally, both BP and BP-OSD exhibit an error floor at low physical error rates. Upon examining the decoding failures, we find that many are due to low-weight (e.g., weight-3) errors that fall into trapping sets. Our BP-SF algorithm circumvents this issue by decoding flipped syndromes, which effectively reduces the number of errors in decoding attempts.
@@
Figure 5: Error rates of the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the code capacity model.
@@
Another example is the $\llbracket 288,12,18\rrbracket$ BB code from [2]. As shown in Fig. 6, our BP-SF decoder performs similar to the BP-OSD decoder while using fewer than $50\times(20+1)=1050$ iterations per decoding attempt and the latency is still 100 iterations considering full parallelization, as we analyzed above.
@@
Figure 6: Error rates of the $\llbracket 288,12,18\rrbracket$ BB code under the code capacity model.
V-B Circuit-Level Noise Model
@@
Under the circuit-level noise model, errors are injected uniformly across gates and measurements. They can propagate through the circuit as it runs. We use Stim [10] to generate the syndrome extraction circuit, the detector error model, and the corresponding parity-check matrix. In this matrix, each row represents a detector event, and each column corresponds to a specific error mechanism. Following the convention in prior literature, we perform $d$ rounds of syndrome extraction and define the logical error rate per round as
$\text{LER Per Round}=1-(1-\text{LER})^{\frac{1}{d}},$ (11)
@@
In this model, the number of error mechanisms is typically much larger than the number of qubits, resulting in very large parity-check matrices. Consequently, flipping a single bit as in the code capacity model is often insufficient for the BP decoder to converge. On the other hand, exhaustively decoding all trial vectors with weight up to a threshold is computationally expensive. To address this, we adopt a sampling-based approach. Given a maximum trial vector weight $w_{max}$, we randomly sample $n_{s}$ trial vectors for each weight in $\{1,\dots,w_{max}\}$, resulting in a total of $n_{s}\times w_{max}$ trial vectors per failed BP decoding attempt.
@@
Figure 7 shows the logical error rates of different decoders on the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model. In this setting, the number of error mechanisms is approximately 8,000, which is significantly larger than the code length used in the code capacity model. Therefore, we expand the candidate set size in our BP-SF decoder. Specifically, our decoder uses up to 3,100 BP iterations (based on $w_{max}=6,n_{s}=5$) and achieves logical error rates that are slightly higher but still comparable to that of the BP-OSD decoder, which uses a maximum of 1,000 iterations and OSD order 10.
@@
Figure 7: Error rates of the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model.
Fig. 8 shows the logical error rates (LER) of different decoders on the $\llbracket 288,12,18\rrbracket$ BB code under the circuit-level noise model. The proposed BP-SF decoder exhibits a slightly higher LER than the BP1000–OSD10 decoder. We note that all decoders for this code employ the layered BP variant, as the regular BP shows significantly worse LER performance. This behavior is likely caused by symmetric trapping sets, as suggested in [20], which can lead to large variations in LER depending on the BP scheduling strategy.
@@
Figure 8: Error rates of the $\llbracket 288,12,18\rrbracket$ BB code under the circuit-level noise model. Layered BP decoding is used for all decoders except the one shown with a dashed line, which is BP-SF using flooding BP.
@@
Fig. 9 shows the logical error rates of different decoders on the $\llbracket 154,6,16\rrbracket$ BB code under the circuit-level noise model. Our BP-SF decoder uses up to 6,000 iterations (based on $w_{max}=6,n_{s}=10$. However, since all trial syndromes are decoded in parallel, the effective decoding latency corresponds to only 200 iterations if the attempts are run in parallel. Our BP-SF decoder achieves logical error rates that are slightly higher but still comparable to that of the BP-OSD decoder at lower physical error rate, which uses a maximum of 1,000 iterations and OSD order 10. In the higher physical error rate regime, our BP-SF decoder exhibits logical error rates that are higher than those of BP-OSD but still consistently lower than baseline BP decoding.
@@
Figure 9: Error rates of the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the circuit-level noise model.
@@
Fig. 10 shows the logical error rates of different decoders on the $\llbracket 126,12,10\rrbracket$ coprime-BB code under the circuit-level noise model. Our BP-SF decoder uses up to approximately 3,000 BP iterations to achieve a logical error rate comparable to that of the BP1000-OSD10 decoder. By increasing both $n_{s}$ and $w_{\max}$, we are able to further reduce the logical error rate to slightly below that of the BP-OSD decoder. However, this improvement comes at the cost of increased complexity, requiring up to 10,000 BP iterations.
Figure 10: Error rates of the $\llbracket 126,12,10\rrbracket$ coprime-BB code under the circuit-level noise model.
To demonstrate that our decoder generalizes across different classes of qLDPC codes, we evaluated BP-SF on the $\llbracket 225,16,8\rrbracket$ subsystem hypergraph product simplex (SHYPS) code [16], as shown in Fig. 11. The proposed BP-SF achieves nearly an identical logical error rate (LER) performance to the BP1000–OSD10 decoder, using $w_{\text{max}}$ and only $n_{s}=5$, i.e., with fewer parallel trials than required for other codes.
Figure 11: Error rates of the $\llbracket 225,16,8\rrbracket$ SHYPS code under the circuit-level noise model.
V-C Complexity and Parameter Selection